If 14L of carbon disulfide (density the same as 237 K) reacts with 40g of oxygen to form sulfur dioxide and carbon dioxide at 180K, then what is the final pressure of the 100L container? (O2 and SO2 have Van deer Waal constants of a = 1.382 & 6.803 L^2 bar/mol^2 and b = 0.03186 & 0.05636 L/mol respectively)
1
Expert's answer
2020-12-08T05:30:47-0500

"P=P1+P2"


"P1=\frac{nRT}{V-nb}-n^2\frac{a}{V^2}= \frac{0.625\times8.31\times180}{100-0.625\times0,00003186005636}-0.625^2(\frac{0.13826803}{100^2})=9350.2317112562 Pa"

"P2=\frac{nRT}{V-nb}-n^2\frac{a}{V^2}= \frac{0.625\times8.31\times180}{100-0.625\times0,00003186005636}-0.625^2(\frac{0.13826803}{100^2})=1732303.2333654 Pa"

"P=P1+P2=9350.2317112562Pa+1732303.2333654 Pa=1 741 653.47Pa=1741.65kPa"